QUESTION 9 Please show all work. For full credit provide equations used. Please use correct significant figures and correct utemp is : 298K
Answer #1

Reaction:
2Fe3+(aq) + Hg22+ (aq) \rightleftharpoons 2Fe2+ (aq) + 2Hg2+ (aq) Kc = 9.1 x 10-6 at 298 K

a.

This is the standard temperature, so:
\Delta G^{o} = -RTlnK

Here, R = 8.314 J/mol-K , T = 298 K , K = 9.1 x 10-6  

\Delta G^{o} = - 8.314 J/mol.K \times 298K \times ln(9.1\times 10^{-6})

\Delta G^{o} = - 2477.57 J/mol \times[ ln(9.1) + ln (10^{-6})]

AG° = -2477.57J/mol x [2.2 – 6(in 10)

\Delta G^{o} = - 2477.57 J/mol \times[ 2.2 -6 (2.3)]

\Delta G^{o} = - 2477.57 J/mol \times[ 2.2 -13.8]

\Delta G^{o} = - 2477.57 J/mol \times[-11.6]

\Delta G^{o} = +28739.8 J/mol

\Delta G^{o} = +28. 7 kJ/mol

b.

Q for the given reaction can be written as:

Q = \frac{ [Fe^{2+}]^{2} [Hg^{2+}]^{2}}{[Fe^{3+}]^{2} [Hg_{2}^{2+}]}

If all concentrations are 1 . Then, Q = 1 and as we know Kc = 9.1 x 10-6 .

Hence Q > K -> so reaction will move towards left (backward direction) to produce more of the reactants.

c.

\Delta G = \Delta G^{o} +RTlnQ

Here \Delta G^{o} = +28. 7 kJ/mol = 28700 J/mol , R = 8.314 J/mol.K , T = 298K  
and, Q can be calculated using given concentrations: [Fe3+] =0.20M [Hg22+] =0.010M [Fe2+] = 0.010M [Hg2+] = 0.025M

Q = \frac{ [Fe^{2+}]^{2} [Hg^{2+}]^{2}}{[Fe^{3+}]^{2} [Hg_{2}^{2+}]} = \frac{(0.010)^{2} (0.025)^{2}}{(0.20)^{2}(0.010)}

Q = \frac{10^{-4} \times 6.25 \times 10^{-4}}{4 \times 10^{-2}\times 10^{-2} }

Q = 1.56 \times 10^{-4}

Here, Q > K -> so reaction will move towards left (backward direction) to produce more of the reactants.

Now,

\Delta G = \Delta G^{o} +RTlnQ

\Delta G = 28700J/mol + 8.314 J/mol.K \times 298K \times ln(1.56 \times 10^{-4})

\Delta G = 28700J/mol + 2477.6 J/mol \times[ ln(1.56) + ln(10^{-4})]

\Delta G = 28700J/mol + 2477.6 J/mol \times[ 0.44 -4 \times ln(10)]

\Delta G = 28700 + 2477.6 \times[ 0.44 -4 \times 2.3]J/mol

\Delta G = 28700 + 2477.6 \times[ 0.44 -9.2]J/mol

\Delta G = 28700 + 2477.6 \times[-8.76]J/mol

\Delta G = 28700 -21703.8 \, \, J/mol

\Delta G = 6996.2 \, \, J/mol

\Delta G = 6. 99 \, \, kJ/mol